### v1

# GraspXL: Generating Grasping Motions for Diverse Objects at Scale

Abstract
Human hands possess the dexterity to interact with diverse objects such as grasping specific parts of the objects and/or approaching them from desired directions. More importantly, humans can grasp objects of any shape without object-specific skills. Recent works synthesize grasping motions following single objectives such as a desired approach heading direction or a grasping area. Moreover, they usually rely on expensive 3D hand-object data during training and inference, which limits their capability to synthesize grasping motions for unseen objects at scale. In this paper, we unify the generation of hand-object grasping motions across multiple motion objectives, diverse object shapes and dexterous hand morphologies in a policy learning framework GraspXL. The objectives are composed of the graspable area, heading direction during approach, wrist rotation, and hand position. Without requiring any 3D hand-object interaction data, our policy trained with 58 objects can robustly synthesize diverse grasping motions for more than 500k unseen objects with a success rate of 82.2%. At the same time, the policy adheres to objectives, which enables the generation of diverse grasps per object. Moreover, we show that our framework can be deployed to different dexterous hands and work with reconstructed or generated objects. We quantitatively and qualitatively evaluate our method to show the efficacy of our approach. Our model and code will be available. Project page: https://eth-ait.github.io/GraspXL.
Keywords:
Motion synthesis Hand-object interaction Dexterous manipulation
Figure 1: Large-scale Grasping Synthesis. Our method, GraspXL, can be used to generate large-scale grasps with robotic hands, and the MANO hand model. Here we show large-scale generated results, better viewed when zoomed in.
Figure 2: Objective-driven Grasping Motion Synthesis. Given a hand model and an object, our goal is to synthesize grasp motions that adhere to high-level objectives, which may consist of one or multiple objectives including graspable areas (indicated by the shadow), heading directions (indicated by the red arrow), wrist rotations (indicated by the black arrow), and positions of the hand (indicated by the green dot). For each sequence, darker hand represents more recent in time.
1 Introduction
In our daily lives, we constantly engage with a wide variety of objects, from taking an apple out of a bowl and handling a knife by its grip to lifting a pair of headphones off the floor. This routine showcases the remarkable dexterity and adaptability of human hands, which effortlessly achieve complex tasks such as precisely grasping objects of different shapes in specific areas and from certain directions. Impressively, humans achieve this without needing object-specific pre-training, enabling us to manipulate items of any shape with ease. Therefore, the ability to generate versatile grasping motions that adhere to certain motion objectives – like precise graspable areas and specific heading directions – holds significant benefits for fields like animation and robotic grasping [20, 7, 1].
In this paper, we present GraspXL, a policy learning framework capable of generating motions for a wide variety of objects, motion objectives, and hand morphologies, which is shown as Fig. 2. Our method does not rely on any 3D hand-object data to train but can robustly generalize to grasp a broad range of unseen objects. Consequently, our approach significantly scales hand-object motion generation, accommodating over half a million unseen objects, and we show some examples in Fig. 1. Existing methods for generating hand-object motions struggle with scalability to unseen objects due to their dependence on pose references [51, 9], their need for time-intensive optimization [45, 7], or their limitation to objects encountered during training [27]. Furthermore, these existing approaches cannot, when applied off-the-shelf, generate interacting motions that fulfill multiple motion objectives.
Scaling objective-driven grasping motion generation to a wide variety of unseen objects poses several challenges. First, there is the generalization ability, where the learning framework should be general enough to handle different object shapes, dexterous hand models, and motion objectives. It is necessary to design a framework that avoids specific assumptions about the object shape and hand model. Second, the model must establish stable grasping while adhering to multiple objectives. Given the variety of object shapes and objectives, the model needs to learn stable grasping while satisfying multiple objectives. However, these high-level goals may negatively affect each other during training as the exploration of objectives can lead the model into local optima. In such cases, the objectives are followed initially, but no stable grasp is reached due to the object movement caused by contact, making the learning difficult and increasing the requirements for control precision.
We formulate GraspXL in the reinforcement learning paradigm and leverage physics simulation. To allow a policy to react to varying object shapes, we capture the general shape features of diverse objects with the vectors from each finger joint to the nearest point on the object surface. To handle multiple objectives, we introduce a control scheme, dubbed objective-driven guidance, that guides the hand towards the desired objective(s). To achieve generalization across diverse hands, we propose a general reward function composed of a grasping reward term and an objective reward term that is agnostic to the hand morphology. Finally, to tackle the difficulty of learning stable grasping while satisfying the target objectives, we propose a learning curriculum to decompose the learning process to objective learning and grasp learning. Specifically, we first train the policy on stationary objects with a larger objective reward to learn precise finger motions for the objectives. We then fine-tune the policy on non-stationary objects with a larger grasping reward to promote stable grasping.
In our experiments, we evaluate methods for multi-objective grasping motion synthesis on PartNet [29] and ShapeNet [4]. We enhance SynH2R [7], the only method offering controllability in hand heading direction, by incorporating more detailed motion objectives. Our approach, unlike SynH2R [7], achieves higher performance without time-intensive optimization for reference poses, yielding a 30% increase in success rates and reducing objective errors by 30%-50%. Additionally, we demonstrate our method’s broad applicability and generalization across a diverse range of conditions: it effectively handles over half a million objects from large-scale 3D datasets [12], adapts to objects from text-to-3D generation methods [30], applies to objects from 3D reconstruction techniques [17], and operates with various robotic hands, including Shadow [36], Allegro [44], and Faive [41]. We validate our method’s superiority over others through quantitative and qualitative measures and highlight its broad generalization capabilities. Additionally, we dissect our framework’s critical elements and investigate the impact of various objective combinations on performance.
In summary, our contributions are: 1) GraspXL, a framework that synthesizes grasping motions on a large scale (500k+) of unseen objects, without relying on hand-object datasets during training. 2) A learning curriculum and objective-driven guidance to enable our method to achieve stable grasping while satisfying multiple objectives. 3) We show that our method is general enough to be deployed on reconstructed or generated objects and different dexterous hand platforms. We will release the code and models upon publication.
2 Related Work
Table 1: Comparison with existing grasping motion synthesis methods.
Method Multiple Different Data-agnostic Number of Novel
Objectives Hands Inference Test Objects
DexVIP [27] $\times$ $\times$ $\times$ 0
D-Grasp [9] $\times$ $\times$ $\times$ 3
UniDexGrasp [45] $\times$ $\times$ 100
UniDexGrasp++ [43] $\times$ $\times$ 100
SynH2R [7] $\times$ $\times$ 1,174
GraspXL (Ours) 503,409
We categorize related works into hand-object interaction synthesis and dexterous robot hand manipulation. Table 1 compares different grasping motion synthesis methods with ours regarding the use of objectives, different hand morphologies, whether they require datapoints at inference time, and the number of reported results on unseen objects.
2.1 Hand-object Interaction Synthesis
In the literature, hand interaction primarily focuses on hand(-object) reconstruction [40, 19, 26, 3, 46, 18, 14, 53, 16], static grasp synthesis [11, 42, 23, 48] and temporal hand-object motion synthesis [52, 9, 8, 2, 38, 20, 50]. This work concentrates on the latter. In synthesis, some previous methods use pure data-driven approaches to synthesize human hand manipulation sequences and rely on post-processing to enhance physical plausibility [52, 50]. Data-driven methods are supervised by 3D hand-object annotated sequences during training [52, 20, 38, 8], and some methods rely on references like wrist trajectories [50] during inference. However, acquiring accurate 3D hand-object data is infeasible to scale because of expensive capture setups [19, 39]. Due to limited training data, data-driven methods are constrained by their training distribution, making generalization challenging, especially for conditional generation tasks with motion objectives. Moreover, their dependence on data restricts their applicability across different dexterous hand platforms such as Shadow Hand [36] and Allegro Hand [44].
Some works leverage physics simulation within reinforcement-learning frameworks to alleviate the data requirement and ensure physical feasibility. Christen et al. [9] generate natural grasp sequences from captured or reconstructed static grasp references. Zhang et al. [51] achieve two-hand grasp and articulation with a single frame reference grasp. However, both methods lack extensive evaluation to demonstrate their generalization ability, and their reliance on references restricts them from scaling to more objects. Some methods [7, 45] first generate grasp reference poses and then synthesize grasping motions accordingly for thousands of objects. However, they either rely on a time-consuming optimization process [7, 45] or are limited by the diversity in the pre-collected dataset [45]. Furthermore, the pre-defined references may not be physically feasible, which introduces extra disturbances to the generated motions as the hands should adhere strictly to the imperfect references. In contrast to existing approaches, our method does not require grasping references and offers real-time inference capabilities for grasping a wide range of objects, while at the same time enabling the control over multiple motion objectives such as graspable areas, heading directions, wrist rotations, and hand positions.
2.2 Dexterous Robot Hand Manipulation
Dexterous manipulation plays a crucial role in enhancing robot capabilities [49, 37, 47, 24]. With the motivation to learn from humans, some approaches use imitation learning. However, this requires full human demonstrations for both training and inference [32, 6, 25], which are usually expensive to collect. Some other methods utilize retargeted human demonstrations during training [47], use teleoperated sequences as training data [33], or learn a parameterized reward function based on demonstrations [10]. However, the reliance on expensive full grasping sequences during training still limits their generalization ability for out-of-distribution settings. Instead of relying on full trajectories, Xu et al. [45] use static pose references to guide the motion by first predicting the contact map and then optimizing the poses accordingly, with physics-based heuristics used to filter out invalid poses. This process is expensive and makes real-time inference infeasible. In contrast, our method generate motions at real-time while adhere to fine-grained motion objectives. Furthermore, without reliance on hand-object interaction data, our method can easily scale to 500k objects.
Another line of research learns grasping with RL without conditioning on reference grasps [5, 13, 31]. Wan et al. [43] propose a curriculum learning framework to distill a state-based policy into a vision-based policy. They generalize to many objects and achieve impressive results on vision-based grasping, whereas we focus on generalizable grasping with different objectives in the state-based setting. Similar to ours, some recent methods can achieve affordance-aware grasps. Mandikal et al. [28] train an affordance-aware grasp policy with RL and evaluated with only 24 novel objects, and further introduce a hand pose prior from YouTube videos for natural hand configuration [27] which is only evaluated with objects used for training. Agarwal et al. [1] learn a category-level policy that grasps the objects by affordance areas, but they are limited to grasp their training categories and can generalize to only 1 unseen category. Moreover, all of these methods only generate a single pose per object. In contrast, our method can generate diverse affordance-aware grasping poses that can be controlled via motion objectives, and can be deployed for over 500k unseen objects.
Figure 3: Overview of GraspXL. As shown in the top row, our method can utilize captured, generated, or reconstructed objects, and different dexterous hand platforms such as MANO, Shadow, Allegro or Faive. With given object and hand model, the policy takes different objectives and states as inputs (on the left), and outputs dynamic grasp motions according to the specific objectives (on the right, accordingly, where darker hand represents more recent in time). The objectives can be the heading direction, wrist rotation, hand position or graspable area, and the states contain the hand state, contact, force and distance of each link with the object, and the object point cloud.
3 GraspXL
Task Definition: As illustrated in Fig. 4, we assume a hand model $h$ with $L$ links where $\mathbf{h}_{i}\in\rm I\!R^{3}$ is the position of the i-th link. The hand pose consists of joint angles $\mathbf{q}\in\rm I\!R^{L\times 3}$ and the global orientation represented by the heading direction $\mathbf{v}\in\rm I\!R^{3}$ and wrist rotation $\omega\in\rm I\!R$ about the directional vector $\mathbf{v}$. As the hand moves, it has linear and angular velocities $\mathbf{u}_{h}\in\rm I\!R^{6}$. We define $\mathbf{m\in\rm I\!R^{3}}$ as the midpoint between the thumb tip and the third joint of the middle finger. We also assume a rigid object $o$ which has a point cloud of 3D vertices $\{\mathbf{o}_{j}\}$. A user may partition the point cloud into graspable/non-graspable points. Specifically, $\{\mathbf{o}_{j}\}=\{\mathbf{o}^{+}_{j}\}\cup\{\mathbf{o}^{-}_{j}\}$ where we split the points into two disjoint sets to specify whether the point should be encouraged to be in contact with the hand or not when motion is generated. As the object moves, it also has linear and angular velocities $\mathbf{u}_{o}\in\rm I\!R^{6}$. The hand links may be in contact $\mathbf{c}\in\{0,1\}^{L}$ with the object using forces in magnitude $\mathbf{f}\in\rm I\!R^{L}$. In particular, the links may contact with the graspable/non-graspable area, which are denoted as $\mathbf{c}^{+}\in\{0,1\}^{L}$ an $\mathbf{c}^{-}\in\{0,1\}^{L}$. Similarly, the force magnitude vector $\mathbf{f}$ can be decomposed into $\mathbf{f}^{+}\in\rm I\!R^{L}$ and $\mathbf{f}^{-}\in\rm I\!R^{L}$.
A user may specify motion objectives $\mathcal{T}$ to define the target heading direction $\bar{\mathbf{v}}$, wrist rotation $\bar{\omega}$, midpoint position $\bar{\mathbf{m}}$, the partition of graspable/non-graspable object point cloud $\{\mathbf{o}^{+}_{j}\}\cup\{\mathbf{o}^{-}_{j}\}$. Given this specification, our goal is to generate a motion sequence that approaches and grasps an object without dropping whilst adhering to the objectives. Note that only the target heading direction $\bar{\mathbf{v}}$ is mandatory to specify in $\mathcal{T}$. If not specified, $\{\mathbf{o}_{j}\}$ will equal to $\{\mathbf{o}^{+}_{j}\}$, $\bar{\mathbf{m}}$ will be the mean of $\{\mathbf{o}_{j}\}$, $\bar{\omega}$ will be zero. See SupMat for more details.
Figure 4: Task definition. (a) The local coordinate system of the hand where the x-axis is the heading direction $\mathbf{v}$, the origin is the midpoint position $\mathbf{m}$ (see text for definition), the rotation about $\mathbf{v}$ is $\omega$. (b) Given an object with user-specified graspable $\{\mathbf{o}^{+}_{j}\}$ and non-graspable vertices $\{\mathbf{o}^{-}_{j}\}$ (labelled in red and blue), the goal of the agent is to approach and grasp the object while satisfying motion objectives $\bar{\mathbf{v}}$, $\bar{\mathbf{m}}$, $\bar{\omega}$, and contact with the graspable area $\{\mathbf{o}^{+}_{j}\}$.
Overview: Figure 3 outlines our RL-based method, GraspXL. Given the objectives $\mathcal{T}$ and the state s obtained from the physics simulation, we utilize a feature extraction layer $\Phi$ to derive the features. Subsequently, the policy $\boldsymbol{\pi}$ takes features as input and generates the actions a, representing the PD-control targets used to compute the torques $\boldsymbol{\tau}$. These torques are then applied to the joints of the hand model in the physics simulation to update the state s, which is subsequently fed back into our feature extraction layer for the next iteration.
Reinforcement Learning Background: The task is formulated as a standard Markov Decision Process (MDP). The goal is to determine the policy $\boldsymbol{\pi}$ that maximizes the expected reward $\mathbb{E}_{\xi\sim\boldsymbol{\pi}}\left[\sum_{t=0}^{T}\mathcal{\gamma}^{t}r_{t}\right]$, where $\mathcal{\gamma}\in[0,1]$ denotes the discount factor, $r_{t}$ represents the reward at time step $t$, and $\mathcal{\xi}=[(\mathbf{s}_{0},\mathbf{a}_{0}),\cdots,(\mathbf{s}_{T},\mathbf{a}_{T})]$ denotes a state-action pair trajectory generated by the policy $\boldsymbol{\pi}$ interacting with the physics simulation. The trajectory $\mathcal{\xi}$ is determined by the transition function $p(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})$, which is governed by the physics simulation along with an initial state distribution $p(s_{0})$. The distribution of a trajectory $\mathcal{\xi}$ is defined as $p_{\theta}(\mathbf{\xi})=p(\mathbf{s}_{0})\prod_{t=0}^{T}p(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t})\boldsymbol{\pi}(\mathbf{a}_{t}|\Phi(\textbf{s}_{t},\mathcal{T}))$. Here, $\boldsymbol{\pi}$ is represented by a neural network.
Feature Extraction: Given a hand-object state s and the motion objectives $\mathcal{T}$, we extract the following features with a conversion function $\Phi(\textbf{s},\mathcal{T})$:
$\Phi(\textbf{s},\mathcal{T})=(\textbf{q},\textbf{d},\mathbf{u}_{h},\mathbf{u}_{o},\textbf{c},\textbf{f},\tilde{\textbf{v}},\tilde{\textbf{m}},\tilde{\omega},\textbf{l}),$ (1)
which includes finger joint angles q, the finger joint tracking error (compared with the target angles for PD controller) d, the hand and object velocities $\mathbf{u}_{h}$ and $\mathbf{u}_{o}$, the contact vector c, the force magnitude vector f. Finally, $\tilde{\textbf{v}},\tilde{\textbf{m}},\tilde{\omega}$ represent the differences between current and target heading directions, midpoint positions, and wrist rotation angles. To represent shape features of an object and help the hand get aware of the graspable/non-graspable areas, we construct distance features $\textbf{l}^{+}\in\rm I\!R^{L\times 3}$ where each row is the difference vector between a link position $\mathbf{h}_{i}$ to the closet object vertex in the graspable part $\{\mathbf{o}^{+}_{j}\}$. Similarly, we construct $\textbf{l}^{-}\in\rm I\!R^{L\times 3}$ for the non-graspable part $\{\mathbf{o}^{-}_{j}\}$. We ablate this component in Section 4.4.
Reward Function: The reward function should guide our policy to learn a solution that can grasp the objects with the desired objectives while at the same time achieving successful grasps. Furthermore, it should be formulated without specific assumptions about the hand morphology so as to be applied on different hand models. As a result, we define our reward function as follows:
$r=r_{\text{goal}}+r_{\text{grasp}},$ (2)
where $r_{\text{goal}}$ is for motion objectives and $r_{\text{grasp}}$ is for successful grasping.
In particular, the motion objective reward $r_{\text{goal}}$ is formulated as:
$r_{\text{goal}}=r_{\text{dis}}+r_{\mathbf{v}}+r_{\omega}+r_{\mathbf{m}}$ (3)
where $r_{\text{dis}}$ promotes approaching the target while avoiding non-graspable areas, and $r_{\mathbf{v}}$, $r_{\omega}$, and $r_{\mathbf{m}}$ reward aligning heading direction, wrist rotation, and midpoint, respectively. Concretely, the term $r_{\text{dis}}$ is defined as
$r_{\text{dis}}=-\sum_{i=1}^{L}\left[w_{d}^{+}(i)||\textbf{h}_{i}-\textbf{o}^{+}_{i}||^{2}-w_{d}^{-}(i)||\textbf{h}_{i}-\textbf{o}^{-}_{i}||^{2}\right]$,
with the weights $w_{d}^{+}(i)\in\rm I\!R$ and $w_{d}^{-}(i)\in\rm I\!R$, the i-th link position ($\textbf{h}_{i}$), the closest graspable/non-graspable object points to the i-th link ($\textbf{o}^{+}_{i}$ and $\textbf{o}^{-}_{i}$). The rewards $r_{\mathbf{v}}$, $r_{\omega}$, and $r_{\mathbf{m}}$ penalize discrepancies in the heading direction v, the wrist rotation angle $\omega$, and midpoint position m from the targets:
$r_{\mathbf{v}}=-w_{\mathbf{v}}||\textbf{v}-\overline{\textbf{v}}||^{2},\ r_{\omega}=-w_{\omega}||\omega-\overline{\omega}||^{2},\ r_{\mathbf{m}}=-w_{\mathbf{m}}||\textbf{m}-\overline{\textbf{m}}||^{2}$.
The grasp reward $r_{\text{grasp}}$ promotes proper contact and natural poses by combining multiple factors:
$r_{\text{grasp}}=r_{\mathbf{c}}+r_{\mathbf{f}}+r_{\text{anatomy}}+r_{\text{reg}}$ (6)
where $r_{\mathbf{c}}$ assesses the contact between fingers and the object, considering both the target and non-target areas ($r_{\mathbf{c}}=w_{c}^{+}\left\lVert\textbf{c}^{+}\right\rVert^{2}-w_{c}^{-}\left\lVert\textbf{c}^{-}\right\rVert^{2}$). The force term, $r_{\mathbf{f}}$, encourages the contact forces with graspable area and punishes the contact forces with the non-graspable area, capped by a factor proportional to the object’s weight $w_{o}$: ($r_{\mathbf{f}}=w_{f}^{+}\textbf{c}^{+}\cdot min(\textbf{f}^{+},\lambda w_{o})-w_{f}^{-}\textbf{c}^{-}\cdot min(\textbf{f}^{-},\lambda w_{o})$). $r_{\text{reg}}$ penalizes excessive velocities to ensure stability ($r_{\text{reg}}=-w_{h}||\mathbf{u}_{h}||^{2}-w_{o}||\mathbf{u}_{o}||^{2}$). If the hand model is MANO [34], we additionally apply the anatomy reward $r_{\text{anatomy}}$ [46] on joint angles $\mathbf{q}$ for natural poses.
Curriculum: Generating grasping motion with objectives requires a policy that establishes stable grasps while achieving specific goals. This is complicated because of the potential for adverse outcomes like object flipping caused by contact when trying to accomplish the objectives. To mitigate this, we introduce a learning curriculum: we start by training the policy on static objects with increased $r_{\text{goal}}$ to hone precise finger movements for objectives. Training progresses to moving objects with a higher $r_{\text{grasp}}$, enhancing wrist movements for secure, non-slip grasps. The effectiveness of this approach is discussed in Section 4.4.
Objective-driven hand guidance: Diverse objectives complicate policy exploration due to the need for varied wrist movements. To address this, we present a simple-yet-effective method to guide the hand during training and inference, improving exploration and control precision. Essentially, we compute the differences between the target and the current values for the heading direction $\mathbf{v}$, the wrist rotation angle $\omega$, and the midpoint position $\mathbf{m}$. These differences are then applied directly as bias terms for the wrist’s 6 degrees of freedom (DoF) PD-controller to guide the wrist toward motion objectives. This simple trick promotes quicker convergence and boosts performance, as detailed in Section 4.4.
4 Experiments
This section involves several experiments to evaluate our method’s effectiveness and generalization capabilities. We detail the experimental setups in Section 4.1, compare our method’s performance against others in Section 4.2, and examine its generalization across various unseen objects and hand models in Section 4.3. Lastly, we ablate our method’s components and analyze the influence of different objective combinations in Section 4.4.
4.1 Experimental Setup
Datasets: To construct the training set, we randomly select 26 objects and 32 objects from ShapeNet [4] and PartNet [29] respectively. To demonstrate our method’s generalization to large-scale object datasets, we use Objaverse [12]. Since not all objects in the three datasets are suitable for rigid-body grasping (e.g. a piece of paper), we filter out objects with such shapes. After filtering and preprocessing, we have 48, 3993 and 503k objects in PartNet, ShapeNet, and Objaverse for test, respectively. The PartNet test set contains unseen objects from categories seen during training, while the objects of the other two test sets are completely novel. Please see SupMat for details on dataset filtering and preprocessing.
Training Details: We use PPO [35] for RL training with the RaiSim [22] physics simulator. Experiments are conducted using a single Nvidia RTX 6000 GPU with 128 CPU cores. To improve data diversity for better generalization, we sample objectives during training with the following heuristics: we randomly sample the target heading direction $\bar{\mathbf{v}}\in\rm I\!R^{3}$ and wrist rotation angle $\bar{\omega}\in[0,2\pi)$ while ensuring that the graspable part $\{\mathbf{o}^{+}_{j}\}$ of the object is narrower than 12 cm between the thumb and the other fingers (along the y axis of the hand local coordinate in Fig. 4a). As a result, the object will not be too large to be grasped. To find a midpoint, we randomly sample a point on the graspable surface $\{\mathbf{o}^{+}_{j}\}$.
Evaluation Protocol: We measure the stability of the grasps with a success metric metric. To assess the adherence to the objectives, we compute several metrics and utilize the ShapeNet and PartNet datasets, which contains part-based objects. Midpoint Error (Mid. Error): The mean Euclidean distance between the final midpoint position $\mathbf{m}$ and the target $\bar{\mathbf{m}}$ measured in centimeters. Heading Error (Head. Error): The mean geodesic distance between the final heading direction $\mathbf{v}$ and the target $\bar{\mathbf{v}}$ measured in radian. Wrist Rotation Error (Rot. Error): The mean absolute error between the final wrist rotation $\omega$ and the target $\bar{\omega}$ measured in radian. Contact Ratio: The ratio between the number of links contacted with the graspable part $\{\mathbf{o}^{+}_{j}\}$ and the entire object $\{\mathbf{o}_{j}\}$. Grasping Success Rate (Suc. Rate): A grasp is determined as success if the object is lifted higher than 10cm and remains stable without falling until the sequence terminates.
Baselines: We choose SynH2R as main baseline, because it offers some degree of controllability in the motion generation, (in terms of heading direction $\mathbf{v}$), we adapt it with more fine-grained motion objectives. In particular, we have the following baselines. SynH2R: The original method generates grasping motion by first generating static grasping reference poses with an optimization procedure. Then it uses an RL-based policy to approach and follow the grasping reference pose. To adapt this method for objective-driven synthesis beyond heading direction $\mathbf{v}$, we add additional optimization terms to include the wrist rotation error $\omega$ and the midpoint error $\mathbf{m}$. We also encourage the contacts with the graspable part while punishing the contacts with the non-graspable part during the optimization. The RL-based policy stays the same as the original one in SynH2R, which is trained with the generated grasping references. SynH2R-PD: The grasping references are generated the same way as in SynH2R above. We then set the reference poses as targets for the PD controller following [9].
4.2 Method Comparison
[…]
Table 2: Method Comparison on PartNet and ShapeNet.
PartNet Test Set ShapeNet Test Set
Method Suc. Rate Mid. Error Head. Error Rot. Error Contact Ratio Suc. Rate Mid. Error Head. Error Rot. Error
[%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$ [%] $\uparrow$ [%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$
PD 26.5 4.30 0.767 0.857 13.0 21.9 4.60 0.850 0.964
SynH2R 82.3 4.06 0.522 0.568 53.4 65.8 4.49 0.642 0.688
Ours 95.0 2.85 0.270 0.306 86.7 81.0 3.22 0.292 0.338
PartNet Evaluation: We use the PartNet test set to evaluate all methods with the given objectives. For each object, we calculate the average performance among 25 randomly sampled heading directions $\bar{\mathbf{v}}$, wrist rotations $\bar{\omega}$ and midpoints $\bar{\mathbf{m}}$. Table 2 shows that our method significantly outperforms the baselines across all metrics. In particular, our method more closely follows the objectives in terms of hand grasping position (Midpoint Error), heading direction (Heading Error), wrist rotation (Wrist Rotation Error), and contact points with the graspable/non-graspable parts (Contact Ratio). At the same time, we achieve the most stable grasps (Success Rate).
ShapeNet Evaluation: While PartNet specializes in part-based objects at a smaller scale, we extend our evaluation to ShapeNet to test on more diverse and unseen objects. We measure average performance across 5 randomly sampled heading directions $\bar{\mathbf{v}}$, wrist rotations $\bar{\omega}$ and midpoints $\bar{\mathbf{m}}$ per object. As shown in Table 2, our method outperforms all baselines, demonstrating less decrease in performance when comparing PartNet and ShapeNet, thereby proving superior generalization. The Grasp Success Rate has a drop for all the methods (even for the non-learning-based SynH2R-PD baseline), which we claim is caused by some objects that are too large or too heavy to be grasped. It is worth noting that generating grasp pose references for the ShapeNet test set using the baselines requires approximately a week, whereas our method can directly perform real-time inference in simulation, showcasing its efficiency. Please refer to SupMat for further qualitative comparisons.
4.3 Generalization
Generalization to Large-scale Object Dataset: Our method’s scalability was tested using the large Objaverse [12] dataset, resulting in a test set of over half a million objects of three different sizes: small, medium and large (refer to SupMat for preprocessing details). Performance on this set, as detailed in Table 3, is comparable to our ShapeNet results, underscoring our method’s ability to scale. Specifically, medium-sized objects offer optimal grasping success. Smaller objects enable higher success rates and more accurate grasp positioning but suffer from larger heading direction and wrist rotation errors, as smaller objects are easier to grasp but also easier to rotate in hand. Overall, our method consistently shows excellent performance across different object sizes, confirming its effectiveness for varying scales.
Table 3: Generalization to the Large-scale Objaverse Dataset.
Objects Suc. Rate [%] $\uparrow$ Mid. Error [cm] $\downarrow$ Head. Error [rad] $\downarrow$ Rot. Error [rad] $\downarrow$
Small 85.9 3.20 0.311 0.362
Medium 84.5 3.16 0.274 0.315
Large 79.0 3.50 0.271 0.306
Average 82.2 3.32 0.279 0.319
Table 4: Generalization to Generated and Reconstructed Objects.
Objects Suc. Rate [%] $\uparrow$ Mid. Error [cm] $\downarrow$ Head. Error [rad] $\downarrow$ Rot. Error [rad] $\downarrow$
Generated Obj. 88.4 2.85 0.310 0.361
Reconstructed Obj. (in the wild) 77.5 3.68 0.222 0.281
Reconstructed Obj. (YCB) 74.0 3.84 0.207 0.247
Ground-truth Obj. (YCB) 73.7 3.63 0.259 0.301
Generalization to Reconstructed and Generated Objects: Our method not only synthesizes grasping motions for standard 3D assets [12, 29, 4] but also effectively handles reconstructed and generated objects, expanding its usability. We evaluated its performance on objects reconstructed via HOLD [17] and those generated by DreamFusion [30]. We analyzed the average performance for each object across 25 random sets of motion objectives, with findings detailed in Tab. 4. The results are presented in Tab. 4. Despite the objects being novel and containing severe artifacts (see the Rubik’s Cube in Fig. 2), our method maintains similar performance across all metrics compared to other experiments. Notably, the performance of reconstructed HO3D objects is similar with the performance on ground-truth objects. This shows our method’s robustness towards reconstruction noise in object meshes. Please refer to the SupMat for a more detailed explanation of the setting.
Table 5: Generalization to Different Hand Models.
PartNet Test Set ShapeNet Test Set
Hand Suc. Rate Mid. Error Head. Error Rot. Error Contact Ratio Suc. Rate Mid. Error Head. Error Rot. Error
Model [%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$ [%] $\uparrow$ [%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$
Allegro 95.3 4.38 0.291 0.300 81.1 83.4 4.45 0.271 0.292
Shadow 94.0 3.57 0.317 0.320 83.4 83.2 3.67 0.363 0.381
Faive 95.8 2.85 0.228 0.243 88.7 82.4 3.44 0.250 0.262
MANO 95.0 2.85 0.270 0.306 86.7 81.0 3.22 0.292 0.338
Figure 5: Generated Motions of Different Hands with the Same Objectives. We require the hands to approach from right and grasp the upper part of the glass.
Generalization on Different Robotic Hands: We assess our framework’s generalization capabilities on various dexterous robotic hands, including Shadow Hand [36], Allegro Hand [44], and Faive Hand [41], each differing in size and joint structure (Faive Hand has 30, Shadow Hand 22, Allegro Hand 16, and MANO 45 finger joints). Our method adapts seamlessly to different models by adjusting hand state and action space dimensions. According to Tab. 5, all tested hands showed similar success rates. However, the Allegro Hand has greater Mid. Errors and lower Contact Ratios due to its larger size, which makes it difficult to achieve precise position control. The Shadow Hand shows marginally higher objective errors due to its broad, flat palm which limits its dexterity. Despite these structural differences affecting performance metrics, all hands achieved commendable results, highlighting our method’s ability to generalize. This suggests our framework’s capacity to deal with hand models with different morphologies. Figure 5 illustrates the variety of grasping motions generated for different hands under the same motion objectives.
4.4 Ablation and Analysis
Ablation: We conduct an ablation study to evaluate the impact of different components in our method. Specifically, we assessed variations without the hand guidance technique (w/o Guidance), the joint distance features (w/o Distance), and the learning curriculum (w/o Curriculum). Results are detailed in Table 6. The version without the curriculum slightly outperforms in contact ratio on the PartNet set and maintains similar success rates across tests but suffers significantly in objective-related metrics. We also notice the absence of the curriculum makes the results more sensitive to the exact reward coefficients, leading to more grasp failure with slightly higher objective rewards or poor objective precision with higher grasping rewards. These underscore the curriculum’s role in decoupling the learning of stable grasping and objective fulfillment, which can help the policy avoid getting stuck in local optima caused by their influence on each other during training. The model without the joint distance features gets a slightly larger midpoint error and significantly underperforms in all other aspects, highlighting the distance features’ value in understanding object shapes and adjusting to contact-induced movements. Lastly, excluding hand guidance worsens all performance metrics, proving its efficacy in directing the hand to achieve the desired grasp.
Table 6: Effects of Different Components in GraspXL.
PartNet Test Set ShapeNet Test Set
Model Suc. Rate Mid. Error Head. Error Rot. Error Contact Ratio Suc. Rate Mid. Error Head. Error Rot. Error
[%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$ [%] $\uparrow$ [%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$
w/o Guidance 90.0 3.22 0.394 0.425 82.2 68.5 3.74 0.455 0.528
w/o Distance 81.6 2.90 0.419 0.475 84.2 70.7 3.34 0.467 0.510
w/o Curriculum 96.2 4.12 0.381 0.462 88.8 79.6 4.60 0.396 0.461
Ours 95.0 2.85 0.270 0.306 86.7 81.0 3.22 0.292 0.338
Table 7: Generation Performance with Different Objective Combinations.
PartNet Test Set ShapeNet Test Set
Combination Suc. Rate Mid. Error Head. Error Rot. Error Contact Ratio Suc. Rate Mid. Error Head. Error Rot. Error
[%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$ [%] $\uparrow$ [%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$
Direction 96.1 - 0.256 - 90.2 82.5 - 0.268 -
Direction+Rotation 95.1 - 0.263 0.303 87.9 81.8 - 0.276 0.320
Direction+Midpoint 95.2 2.84 0.268 - 88.7 81.4 3.15 0.284 -
Dir.+Rot.+Mid. 95.0 2.85 0.270 0.306 86.7 81.0 3.22 0.292 0.338
Evaluation on Different Objective Combinations: Our method’s adaptability was tested across various objective combinations to gauge performance impacts. Evaluations were conducted for different scenarios: solely controlling heading direction ($\bar{\mathbf{m}}$), adding wrist rotation to heading direction ($\bar{\mathbf{v}}+\bar{\omega}$), combining heading direction with midpoint position ($\bar{\mathbf{v}}+\bar{\mathbf{m}}$), and integrating all three objectives ($\bar{\mathbf{v}}+\bar{\omega}+\bar{\mathbf{m}}$). This variety helps to understand our approach’s efficacy in handling multiple, concurrent objectives and understanding the complexity levels our method can effectively manage. For each objective set, we sample 25 sets of motion objective for PartNet objects and 5 for ShapeNet objects. Results in Tab. 7 show slightly better performance with fewer motion objectives, highlighting the increased control challenge with multiple objectives. This demonstrates our method’s capability to handle varying difficulty levels and its robustness, as indicated by the minimal performance decline.
5 Conclusion
We presented GraspXL, an RL-based method that learns to synthesize grasping motions on large-scale while satisfying one or multiple motion objectives. We introduced a learning curriculum to deal with the control complexity and an objective-driven guidance technique to accelerate exploration during training. To improve the generalization ability, we adopted a joint distance sensor to capture object local shape features. Notably, we demonstrated that our method exhibits a high success rate of 82.2% on a test set with more than 500k unseen objects. Moreover, we showed that our approach works well even when applied to different dexterous hand platforms and with reconstructed or generated objects.
Supplementary Material
The supplementary material contains this document and a video. In Section 6, we provide qualitative results compared against our baseline. We then provide the implementation details about hyperparameters and motion objectives in Section 7. In Section 8, we show the experiment details about data preprocessing and the evaluation with generated and reconstructed objects. Finally, we provide additional experiments in Section 9. We will release code and models.
Figure 6: Qualitative Comparison. SynH2R requires a time-consuming reference generation process, and suffers from noisy references and imperfect reference tracking, which lead to failed grasping or large objective errors.
6 Qualitative Results
We provide qualitative comparisons of our method with SynH2R [7] for objective-driven grasping synthesis in Figure 6. From the figures, we can see that SynH2R either failed to establish a stable grasp due to noisy generated references, or cannot precisely follow the objectives. However, our method does not require a pre-generated reference, and can generate motions with stable grasping while satisfying motion objectives.
7 Implementation Details
7.1 Training Hyperparameters
We use PPO [35] to train our policy and follow the implementation provided in [9]. We present an overview of the important parameters and weight values of the reward function in Table 8 and Table 9.
Table 8: Hyperparameters of GraspXL.
Hyperparameters PPO Value
Epochs 1e4
Steps per epoch 3e4
Environment steps per episode 150
Batch size 2000
Updates per epoch 20
Simulation timestep 2.5e-3s
Simulation steps per action 4
Discount factor $\gamma$ 0.996
GAE parameter $\lambda$ 0.95
Clipping parameter 0.2
Max. gradient norm 0.5
Value loss coefficient 0.5
Entropy coefficient 0.0
Optimizer Adam
Learning rate 5e-4
Hidden units 128
Hidden layers 2
Table 9: Weights of the Reward Function.
Weights Value (1st phase) Value (2nd phase)
$w_{d}^{+}$ 0.3 0.3
$w_{d}^{-}$ 0.06 0.06
$w_{\mathbf{v}}$ 1.0 0.01
$w_{\omega}$ 1.0 0.01
$w_{\mathbf{m}}$ 10.0 10.0
$w_{c}^{+}$ 1.0 1.0
$w_{c}^{-}$ 1.0 1.0
$w_{f}^{+}$ 0.3 0.5
$w_{f}^{-}$ 0.15 0.25
$w_{h}$ 0.001 0.001
$w_{o}$ 0.0 0.1
$w_{anatomy}$ 0.2 0.1
$\lambda$ 5.0 5.0
7.2 Objectives Specification
As explained in Section 3 of the main manuscript, our framework can deal with different combinations of four kinds of objectives: the partition of graspable/non-graspable object point cloud $\{\mathbf{o}^{+}_{j}\}\cup\{\mathbf{o}^{-}_{j}\}$, the heading direction of the hand $\bar{\mathbf{v}}$, the hand wrist rotation $\bar{\omega}$, and the hand midpoint position $\bar{\mathbf{m}}$. We assume the partition $\{\mathbf{o}^{+}_{j}\}\cup\{\mathbf{o}^{-}_{j}\}$ is specified by a user to indicate the desired grasping area, such as a mug handle or a headphone earcup. By default, $\{\mathbf{o}^{+}_{j}\}=\{\mathbf{o}_{j}\}$ and $\{\mathbf{o}^{-}_{j}\}=\emptyset$, which means that the hand can grasp the entire object. $\bar{\mathbf{v}}$ is the only quantity that is mandatory to specify. $\bar{\omega}$ is by default 00 so that the y-axis of the hand local coordinate system (See Fig. 4 in the main manuscript) is parallel to the narrowest edge of the $\{\mathbf{o}_{j}\}$ projection along $\bar{\mathbf{v}}$, which represents the easiest setting for grasping with the given heading direction $\bar{\mathbf{v}}$. $\bar{\mathbf{m}}$ is by default set to be the centroid of $\{\mathbf{o}_{j}\}$.
8 Experimental Details
8.1 Dataset Preprocessing
In order to compare with existing methods and to demonstrate our method’s generalization capabilities, we use the three object datasets: PartNet [29], ShapeNet [4], and Objaverse [12]. Since not all objects are feasible for grasping (such as a piece of paper), we preprocess and filter the objects.
PartNet We select 80 objects from the categories scissors, knife, mug, earphone, and wineglass, which contain part-based segmentation (such as the handle and the main body of a mug). We use the individual parts of each object to specify the graspable/non-graspable areas. Objects are resized to feasible dimensions for grasping.
[…]
8.2 Reconstructed and Generated Objects
We use the eight objects generated with DreamFusion [30] which are available from their project page, and manually scale them to graspable sizes as our test set for reconstructed objects. For reconstructed objects, we use all the fourteen objects reconstructed from HO3D [21] videos and six objects reconstructed from in-the-wild videos reported in HOLD[17]. For comparison, we evaluate with the ground-truth HO3D objects with their original scales. For each object, we randomly sample 25 sets of motion objectives and report the average performance.
9 Additional Experiments
9.1 Training Set Size Effect Evaluation
To show the data efficiency of our method, we train another policy with an enlarged training set composed of 100 PartNet objects (with the graspable area partitions), and 400 ShapeNet objects. We then perform the same evaluation with the PartNet and ShapeNet test sets (See Section 4.1 in the main manuscript). The results are shown in Table 10. Compared with the results with a smaller training set (See Section 4.2 in the main manuscript), there is no significant improvement, which means that a training set composed of 58 objects with diverse shape is sufficient for our framework. This shows the data efficiency of our method. We hypothesize that this is because diverse shapes together with random objectives and initialization during training provide a diverse distribution of grasps and configurations.
Table 10: Comparison with Different Training Set Size.
PartNet Test Set ShapeNet Test Set
Method Suc. Rate Mid. Error Head. Error Rot. Error Contact Ratio Suc. Rate Mid. Error Head. Error Rot. Error
[%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$ [%] $\uparrow$ [%] $\uparrow$ [cm] $\downarrow$ [rad] $\downarrow$ [rad] $\downarrow$
Ours (58 Training Objects) 95.0 2.85 0.270 0.306 86.7 81.0 3.22 0.292 0.338
Ours+ (500 Training Objects) 94.9 3.57 0.307 0.356 88.1 84.3 3.94 0.283 0.346
9.2 SOTA Comparison for Robotic Hand Grasping
To show the effectiveness of our method on robotic applications, we compare our method with two state-of-the-art dexterous robot hand manipulation methods, UniDexGrasp [45] and UniDexGrasp++ [43]. UniDexGrasp samples and optimizes diverse grasp pose references, filters the successful ones with a physics simulator, and then generates diverse grasp motions accordingly. UniDexGrasp++ is specifically trained for higher grasp success rates and generates a single pose for each object. We evaluate our method using their hand model (the Shadow Hand), test sets (one for seen categories and one for unseen categories), and metric (success rate), and compare with their state-based policies. The results are shown in Table 11. Note that both their test sets are unseen to our policy. Our policy is trained with only our 58 training objects instead of their 3000+ training objects. For each object, we randomly sample 25 target approach directions to synthesize diverse grasp motions. Our method outperforms both methods on both test sets. Specifically, although UniDexGrasp only uses filtered references for motion synthesis while ours can directly inference in simulation, our method significantly outperforms it and retains the ability to generate diverse motions. Notably, our method achieves a higher success rate than UniDexGrasp++ on its seen category test set.
Table 11: Comparison with Dexterous Robot Hand Manipulation Methods. We evaluate the performance of our policy using the task setting and metric of UniDexGrasp and UniDexGrasp++ on their test sets. Both ours and UniDexGrasp generate diverse grasps, while UniDexGrasp++ only considers a single grasp for each object.
Method Success Rate [%] $\uparrow$ Success Rate [%] $\uparrow$
(Seen Category) (Unseen Category)
UniDexGrasp (Diverse poses) 71 67
UniDexGrasp++ (Single pose) 84 83
Ours (Diverse poses) 85 89

### v2

@@
 In this paper, we unify the generation of hand-object grasping motions across multiple motion objectives, diverse object shapes and dexterous hand morphologies in a policy learning framework GraspXL.
-The objectives are composed of the graspable area, heading direction during approach, wrist rotation, and hand position.
+The objectives are composed of the graspable area, heading direction, wrist rotation, and hand position.
 Without requiring any 3D hand-object interaction data, our policy trained with 58 objects can robustly synthesize diverse grasping motions for more than 500k unseen objects with a success rate of 82.2%.
@@
 We quantitatively and qualitatively evaluate our method to show the efficacy of our approach.
-Our model and code will be available.
-Project page: https://eth-ait.github.io/GraspXL.
+Our model, code, and the large-scale generated motions are available at https://eth-ait.github.io/graspxl/.
 Keywords:
@@
 Given a hand model and an object, our goal is to synthesize grasp motions that adhere to high-level objectives, which may consist of one or multiple objectives including graspable areas (indicated by the shadow), heading directions (indicated by the red arrow), wrist rotations (indicated by the black arrow), and positions of the hand (indicated by the green dot).
-For each sequence, darker hand represents more recent in time.
+For each sequence, the darker hand represents more recent in time.
+**footnotetext: Now at HKUST(GZ)&HKUST
 1 Introduction
@@
 2) A learning curriculum and objective-driven guidance to enable our method to achieve stable grasping while satisfying multiple objectives.
-3) We show that our method is general enough to be deployed on reconstructed or generated objects and different dexterous hand platforms.
-We will release the code and models upon publication.
+3) A dataset of diverse generated grasp motions for 500k+ objects with different hands.
+4) We show that our method is general enough to be deployed on reconstructed or generated objects and different dexterous hands.
+The code, models, and dataset are released on our project page.
 2 Related Work
@@
 We categorize related works into hand-object interaction synthesis and dexterous robot hand manipulation.
-Table 1 compares different grasping motion synthesis methods with ours regarding the use of objectives, different hand morphologies, whether they require datapoints at inference time, and the number of reported results on unseen objects.
+Tab.
+1 compares different grasping motion synthesis methods with ours regarding the use of objectives, different hand morphologies, whether they require datapoints at inference time, and the number of reported results on unseen objects.
 2.1 Hand-object Interaction Synthesis
-In the literature, hand interaction primarily focuses on hand(-object) reconstruction [40, 19, 26, 3, 46, 18, 14, 53, 16], static grasp synthesis [11, 42, 23, 48] and temporal hand-object motion synthesis [52, 9, 8, 2, 38, 20, 50].
+In the literature, hand interaction primarily focuses on hand(-object) reconstruction [40, 19, 27, 3, 46, 18, 14, 53, 16], static grasp synthesis [11, 42, 24, 48] and temporal hand-object motion synthesis [52, 9, 2, 38, 20, 50].
 This work concentrates on the latter.
 In synthesis, some previous methods use pure data-driven approaches to synthesize human hand manipulation sequences and rely on post-processing to enhance physical plausibility [52, 50].
-Data-driven methods are supervised by 3D hand-object annotated sequences during training [52, 20, 38, 8], and some methods rely on references like wrist trajectories [50] during inference.
+Data-driven methods are supervised by 3D hand-object annotated sequences during training [52, 20, 38], and some methods rely on references like wrist trajectories [50] during inference.
 However, acquiring accurate 3D hand-object data is infeasible to scale because of expensive capture setups [19, 39].
@@
 2.2 Dexterous Robot Hand Manipulation
-Dexterous manipulation plays a crucial role in enhancing robot capabilities [49, 37, 47, 24].
+Dexterous manipulation plays a crucial role in enhancing robot capabilities [49, 47, 25].
 With the motivation to learn from humans, some approaches use imitation learning.
@@
 This process is expensive and makes real-time inference infeasible.
-In contrast, our method generate motions at real-time while adhere to fine-grained motion objectives.
+In contrast, our method generates motions in real-time while adhering to fine-grained motion objectives.
 Furthermore, without reliance on hand-object interaction data, our method can easily scale to 500k objects.
@@
 We also assume a rigid object $o$ which has a point cloud of 3D vertices $\{\mathbf{o}_{j}\}$.
-A user may partition the point cloud into graspable/non-graspable points.
-Specifically, $\{\mathbf{o}_{j}\}=\{\mathbf{o}^{+}_{j}\}\cup\{\mathbf{o}^{-}_{j}\}$ where we split the points into two disjoint sets to specify whether the point should be encouraged to be in contact with the hand or not when motion is generated.
+A user may partition the point cloud into graspable/non-graspable points, specifically $\{\mathbf{o}_{j}\}=\{\mathbf{o}^{+}_{j}\}\cup\{\mathbf{o}^{-}_{j}\}$, where we split the points into two disjoint sets to specify whether the point should be encouraged to be in contact with the hand or not when motion is generated.
 As the object moves, it also has linear and angular velocities $\mathbf{u}_{o}\in\rm I\!R^{6}$.
@@
 (b) Given an object with user-specified graspable $\{\mathbf{o}^{+}_{j}\}$ and non-graspable vertices $\{\mathbf{o}^{-}_{j}\}$ (labelled in red and blue), the goal of the agent is to approach and grasp the object while satisfying motion objectives $\bar{\mathbf{v}}$, $\bar{\mathbf{m}}$, $\bar{\omega}$, and contact with the graspable area $\{\mathbf{o}^{+}_{j}\}$.
-Overview: Figure 3 outlines our RL-based method, GraspXL.
+Overview: Fig.
+3 outlines our RL-based method, GraspXL.
 Given the objectives $\mathcal{T}$ and the state s obtained from the physics simulation, we utilize a feature extraction layer $\Phi$ to derive the features.
@@
 Similarly, we construct $\textbf{l}^{-}\in\rm I\!R^{L\times 3}$ for the non-graspable part $\{\mathbf{o}^{-}_{j}\}$.
-We ablate this component in Section 4.4.
+We ablate this component in Sec.
+4.4.
 Reward Function: The reward function should guide our policy to learn a solution that can grasp the objects with the desired objectives while at the same time achieving successful grasps.
@@
 where $r_{\mathbf{c}}$ assesses the contact between fingers and the object, considering both the target and non-target areas ($r_{\mathbf{c}}=w_{c}^{+}\left\lVert\textbf{c}^{+}\right\rVert^{2}-w_{c}^{-}\left\lVert\textbf{c}^{-}\right\rVert^{2}$).
-The force term, $r_{\mathbf{f}}$, encourages the contact forces with graspable area and punishes the contact forces with the non-graspable area, capped by a factor proportional to the object’s weight $w_{o}$: ($r_{\mathbf{f}}=w_{f}^{+}\textbf{c}^{+}\cdot min(\textbf{f}^{+},\lambda w_{o})-w_{f}^{-}\textbf{c}^{-}\cdot min(\textbf{f}^{-},\lambda w_{o})$).
+The force term, $r_{\mathbf{f}}$, encourages the contact forces with the graspable area and punishes the contact forces with the non-graspable area, capped by a factor proportional to the object’s weight $w_{o}$: ($r_{\mathbf{f}}=w_{f}^{+}\textbf{c}^{+}\cdot min(\textbf{f}^{+},\lambda w_{o})-w_{f}^{-}\textbf{c}^{-}\cdot min(\textbf{f}^{-},\lambda w_{o})$).
 $r_{\text{reg}}$ penalizes excessive velocities to ensure stability ($r_{\text{reg}}=-w_{h}||\mathbf{u}_{h}||^{2}-w_{o}||\mathbf{u}_{o}||^{2}$).
@@
 Training progresses to moving objects with a higher $r_{\text{grasp}}$, enhancing wrist movements for secure, non-slip grasps.
-The effectiveness of this approach is discussed in Section 4.4.
+The effectiveness of this approach is discussed in Sec.
+4.4.
 Objective-driven hand guidance: Diverse objectives complicate policy exploration due to the need for varied wrist movements.
@@
 These differences are then applied directly as bias terms for the wrist’s 6 degrees of freedom (DoF) PD-controller to guide the wrist toward motion objectives.
-This simple trick promotes quicker convergence and boosts performance, as detailed in Section 4.4.
+This simple trick promotes quicker convergence and boosts performance, as detailed in Sec.
+4.4.
 4 Experiments
 This section involves several experiments to evaluate our method’s effectiveness and generalization capabilities.
-We detail the experimental setups in Section 4.1, compare our method’s performance against others in Section 4.2, and examine its generalization across various unseen objects and hand models in Section 4.3.
-Lastly, we ablate our method’s components and analyze the influence of different objective combinations in Section 4.4.
+We detail the experimental setups in Sec.
+4.1, compare our method’s performance against others in Sec.
+4.2, and examine its generalization across various unseen objects and hand models in Sec.
+4.3.
+Lastly, we ablate our method’s components and analyze the influence of different objective combinations in Sec.
+4.4.
 4.1 Experimental Setup
@@
 a piece of paper), we filter out objects with such shapes.
-After filtering and preprocessing, we have 48, 3993 and 503k objects in PartNet, ShapeNet, and Objaverse for test, respectively.
-The PartNet test set contains unseen objects from categories seen during training, while the objects of the other two test sets are completely novel.
+After filtering and preprocessing, we have 48, 3993 and 503k objects in PartNet, ShapeNet, and Objaverse for testing, respectively.
+The PartNet test set contains unseen objects from seen categories, while the objects of the other two test sets are completely novel.
 Please see SupMat for details on dataset filtering and preprocessing.
@@
 Evaluation Protocol: We measure the stability of the grasps with a success metric metric.
-To assess the adherence to the objectives, we compute several metrics and utilize the ShapeNet and PartNet datasets, which contains part-based objects.
+To assess the adherence to the objectives, we compute several metrics and utilize the ShapeNet and PartNet (which contains part-based objects) datasets.
 Midpoint Error (Mid.
@@
 Error): The mean absolute error between the final wrist rotation $\omega$ and the target $\bar{\omega}$ measured in radian.
-Contact Ratio: The ratio between the number of links contacted with the graspable part $\{\mathbf{o}^{+}_{j}\}$ and the entire object $\{\mathbf{o}_{j}\}$.
+textbfContact Ratio: The ratio between the number of links contacted with the graspable part $\{\mathbf{o}^{+}_{j}\}$ and the entire object $\{\mathbf{o}_{j}\}$.
 Grasping Success Rate (Suc.
-Rate): A grasp is determined as success if the object is lifted higher than 10cm and remains stable without falling until the sequence terminates.
-Baselines: We choose SynH2R as main baseline, because it offers some degree of controllability in the motion generation, (in terms of heading direction $\mathbf{v}$), we adapt it with more fine-grained motion objectives.
+Rate): A grasp is determined as a success if the object is lifted higher than 10cm and remains stable without falling until the sequence terminates.
+Baselines: We choose SynH2R as the main baseline, because it offers some degree of controllability in the motion generation, (in terms of heading direction $\mathbf{v}$).
+We adapt it with more fine-grained motion objectives.
 In particular, we have the following baselines.
@@
 For each object, we calculate the average performance among 25 randomly sampled heading directions $\bar{\mathbf{v}}$, wrist rotations $\bar{\omega}$ and midpoints $\bar{\mathbf{m}}$.
-Table 2 shows that our method significantly outperforms the baselines across all metrics.
+Tab.
+2 shows that our method significantly outperforms the baselines across all metrics.
 In particular, our method more closely follows the objectives in terms of hand grasping position (Midpoint Error), heading direction (Heading Error), wrist rotation (Wrist Rotation Error), and contact points with the graspable/non-graspable parts (Contact Ratio).
-At the same time, we achieve the most stable grasps (Success Rate).
+At the same time, we achieve the most stable grasps (Grasping Success Rate).
 ShapeNet Evaluation: While PartNet specializes in part-based objects at a smaller scale, we extend our evaluation to ShapeNet to test on more diverse and unseen objects.
 We measure average performance across 5 randomly sampled heading directions $\bar{\mathbf{v}}$, wrist rotations $\bar{\omega}$ and midpoints $\bar{\mathbf{m}}$ per object.
-As shown in Table 2, our method outperforms all baselines, demonstrating less decrease in performance when comparing PartNet and ShapeNet, thereby proving superior generalization.
-The Grasp Success Rate has a drop for all the methods (even for the non-learning-based SynH2R-PD baseline), which we claim is caused by some objects that are too large or too heavy to be grasped.
+As shown in Tab.
+2, our method outperforms all baselines, demonstrating less decrease in performance when comparing PartNet and ShapeNet, thereby proving superior generalization.
+The Grasping Success Rate has a drop for all the methods (even for the non-learning-based SynH2R-PD baseline), which we claim is caused by some objects that are too large or too heavy to be grasped.
 It is worth noting that generating grasp pose references for the ShapeNet test set using the baselines requires approximately a week, whereas our method can directly perform real-time inference in simulation, showcasing its efficiency.
@@
 Generalization to Large-scale Object Dataset: Our method’s scalability was tested using the large Objaverse [12] dataset, resulting in a test set of over half a million objects of three different sizes: small, medium and large (refer to SupMat for preprocessing details).
-Performance on this set, as detailed in Table 3, is comparable to our ShapeNet results, underscoring our method’s ability to scale.
+Performance on this set, as detailed in Tab.
+3, is comparable to our ShapeNet results, underscoring our method’s ability to scale.
 Specifically, medium-sized objects offer optimal grasping success.
@@
 Overall, our method consistently shows excellent performance across different object sizes, confirming its effectiveness for varying scales.
+The generated grasping motions of different hands are released for further research.
 Table 3: Generalization to the Large-scale Objaverse Dataset.
@@
 Generalization to Reconstructed and Generated Objects: Our method not only synthesizes grasping motions for standard 3D assets [12, 30, 4] but also effectively handles reconstructed and generated objects, expanding its usability.
-We evaluated its performance on objects reconstructed via HOLD [17] and those generated by DreamFusion [30].
+We evaluated its performance on objects reconstructed via HOLD [17] and objects generated by DreamFusion [31].
 We analyzed the average performance for each object across 25 random sets of motion objectives, with findings detailed in Tab.
-4.
-The results are presented in Tab.
 4.
@@
 2), our method maintains similar performance across all metrics compared to other experiments.
-Notably, the performance of reconstructed HO3D objects is similar with the performance on ground-truth objects.
+Notably, the performance of objects reconstructed from HO3D [21] is similar with the performance of ground-truth objects.
 This shows our method’s robustness towards reconstruction noise in object meshes.
-Please refer to the SupMat for a more detailed explanation of the setting.
+Please refer to the SupMat for a more detailed setting explanation.
 Table 5: Generalization to Different Hand Models.
@@
 MANO 95.0 2.85 0.270 0.306 86.7 81.0 3.22 0.292 0.338
-Figure 5: Generated Motions of Different Hands with the Same Objectives. We require the hands to approach from right and grasp the upper part of the glass.
+Figure 5: Generated Motions of Different Hands with the Same Objectives. We require the hands to approach from the right and grasp the upper part of the glass.
 Generalization on Different Robotic Hands: We assess our framework’s generalization capabilities on various dexterous robotic hands, including Shadow Hand [37], Allegro Hand [44], and Faive Hand [41], each differing in size and joint structure (Faive Hand has 30, Shadow Hand 22, Allegro Hand 16, and MANO 45 finger joints).
@@
 This suggests our framework’s capacity to deal with hand models with different morphologies.
-Figure 5 illustrates the variety of grasping motions generated for different hands under the same motion objectives.
+Fig.
+5 illustrates the variety of grasping motions generated for different hands under the same motion objectives.
 4.4 Ablation and Analysis
@@
 Specifically, we assessed variations without the hand guidance technique (w/o Guidance), the joint distance features (w/o Distance), and the learning curriculum (w/o Curriculum).
-Results are detailed in Table 6.
+Results are detailed in Tab.
+6.
 The version without the curriculum slightly outperforms in contact ratio on the PartNet set and maintains similar success rates across tests but suffers significantly in objective-related metrics.
@@
 This variety helps to understand our approach’s efficacy in handling multiple, concurrent objectives and understanding the complexity levels our method can effectively manage.
-For each objective set, we sample 25 sets of motion objective for PartNet objects and 5 for ShapeNet objects.
+For each objective set, we sample 25 sets of motion objectives for PartNet objects and 5 for ShapeNet objects.
 Results in Tab.
@@
 Supplementary Material
-The supplementary material contains this document and a video.
-In Section 6, we provide qualitative results compared against our baseline.
-We then provide the implementation details about hyperparameters and motion objectives in Section 7.
-In Section 8, we show the experiment details about data preprocessing and the evaluation with generated and reconstructed objects.
-Finally, we provide additional experiments in Section 9.
-We will release code and models.
+In Sec.
+6, we provide qualitative results compared against our baseline.
+We then provide the implementation details about hyperparameters and motion objectives in Sec.
+7.
+In Sec.
+8, we show the experiment details about data preprocessing and the evaluation with generated and reconstructed objects.
+Finally, we provide additional experiments in Sec.
+9.
+Our model, code, and the large-scale generated motions are released for future research.
+Check our project page: https://eth-ait.github.io/graspxl/ for more details and visualization.
 Figure 6: Qualitative Comparison.
@@
 6 Qualitative Results
-We provide qualitative comparisons of our method with SynH2R [7] for objective-driven grasping synthesis in Figure 6.
+We provide qualitative comparisons of our method with SynH2R [8] for objective-driven grasping synthesis in Fig.
+6.
 From the figures, we can see that SynH2R either failed to establish a stable grasp due to noisy generated references, or cannot precisely follow the objectives.
@@
 7.1 Training Hyperparameters
-We use PPO [35] to train our policy and follow the implementation provided in [9]. We present an overview of the important parameters and weight values of the reward function in Table 8 and Table 9.
+We use PPO [36] to train our policy and follow the implementation provided in [9]. We present an overview of the important parameters and weight values of the reward function in Tab. 8 and Tab. 9.
 Table 8: Hyperparameters of GraspXL.
@@
 […]
+As the object meshes from the datasets contain no material information for density and friction, we calculate object masses based on the mesh volume for a given fixed density, leading to diverse masses.
+We use the same friction coefficient in simulation for all objects.
 8.2 Reconstructed and Generated Objects
@@
 We then perform the same evaluation with the PartNet and ShapeNet test sets (See Section 4.1 in the main manuscript).
-The results are shown in Table 10.
+The results are shown in Tab.
+10.
 Compared with the results with a smaller training set (See Section 4.2 in the main manuscript), there is no significant improvement, which means that a training set composed of 58 objects with diverse shape is sufficient for our framework.
@@
 Ours+ (500 Training Objects) 94.9 3.57 0.307 0.356 88.1 84.3 3.94 0.283 0.346
-9.2 SOTA Comparison for Robotic Hand Grasping
-To show the effectiveness of our method on robotic applications, we compare our method with two state-of-the-art dexterous robot hand manipulation methods, UniDexGrasp [45] and UniDexGrasp++ [43].
-UniDexGrasp samples and optimizes diverse grasp pose references, filters the successful ones with a physics simulator, and then generates diverse grasp motions accordingly.
-UniDexGrasp++ is specifically trained for higher grasp success rates and generates a single pose for each object.
-We evaluate our method using their hand model (the Shadow Hand), test sets (one for seen categories and one for unseen categories), and metric (success rate), and compare with their state-based policies.
-The results are shown in Table 11.
-Note that both their test sets are unseen to our policy.
-Our policy is trained with only our 58 training objects instead of their 3000+ training objects.
-For each object, we randomly sample 25 target approach directions to synthesize diverse grasp motions.
-Our method outperforms both methods on both test sets.
-Specifically, although UniDexGrasp only uses filtered references for motion synthesis while ours can directly inference in simulation, our method significantly outperforms it and retains the ability to generate diverse motions.
-Notably, our method achieves a higher success rate than UniDexGrasp++ on its seen category test set.
-Table 11: Comparison with Dexterous Robot Hand Manipulation Methods.
-We evaluate the performance of our policy using the task setting and metric of UniDexGrasp and UniDexGrasp++ on their test sets.
-Both ours and UniDexGrasp generate diverse grasps, while UniDexGrasp++ only considers a single grasp for each object.
-Method Success Rate [%] $\uparrow$ Success Rate [%] $\uparrow$
-(Seen Category) (Unseen Category)
-UniDexGrasp (Diverse poses) 71 67
-UniDexGrasp++ (Single pose) 84 83
-Ours (Diverse poses) 85 89
+9.2 Friction Effect Evaluation
+Although we use the same friction coefficient in simulation for all objects, our method can also deal with randomized frictions (±0.3 around the default friction coefficient) as shown in Tab. 11.
+Table 11: Evaluation with random friction coefficients.
+Settings Suc. Rate [%] $\uparrow$ Mid. Error $\downarrow$
+Original experiment 81.6 0.283
+Random friction coefficient 80.5 0.285
+9.3 Realism Evaluation
+To evaluate the realism of our method, we invite 35 participants to score the realism in terms of human likeness, naturalness, smoothness, and hand-object interpenetration.
+They score 10 sets of rendered grasping motions from 1 (worst) to 3 (best), each containing three motions of the same object randomly chosen from ours, HO3D [21], and DexYCB [5].
+The average scores are 2.12, 2.05, and 2.45, respectively.
+Our method has a slightly higher realism score than HO3D as HO3D has some interpenetrations and jitters caused by labeling noise, and DexYCB has the highest score due to accurate annotations.
+9.4 Inference Speed Evaluation
+With a 24-core Intel i9-14900 CPU and an Nvidia RTX 4090 GPU, we generate 100 sequences and calculate the average time consumption per sequence with our method, the baselines SynH2R-PD, and SynH2R [8], leading to 37.15, 37.17, and 0.14 seconds, respectively.
+The most time consumption for SynH2R-PD and SynH2R are caused by the optimization-based static grasping reference pose generation procedure.